# Dual Numbers for Arbitrary Order Automatic Differentiation

### v1

@@ 3 Dual number implementation to arbitrary order
5:     $fr\leftarrow fr+\binom{k}{i}\cdot A\%f(i)\cdot B\%f(k-i)$
@@
7:return $fr$
Algorithm 4 Pseudocode for the function timesd(A,B) which implements the product of the dual numbers A and B
1:$A$: dualzn type, $B$: dualzn type
2:$fr$: dualzn type
3:Allocate $fr\%f(0:\text{order})$
@@
5:     $fr\%f(k)\leftarrow\texttt{timesdzn}(A,B,k)$
@@ 4 The DNAOAD package
With the just presented theory, dual numbers to arbitrary order are implemented in the so-called DNAOAD package. This implementation not only allows for computation of derivatives also to arbitrary order, but also allows the construction of algorithms in the context of dual numbers. Before presenting some interesting applications we first describe the DNAOAD package.
@@ 4.1 Elements of the package
The implementation is available for double (real64) and quadruple (real128) precisions. The user can choose which precision to use. The module for double precision is shown in Listing 2. This module is used by the dualzn_mod module, which is the core of the dual number implementation and contains approximately 2,000 lines of code. The first 28 lines, which also show the functions and operators extended to dual numbers (overloaded), are shown in Listing 3.
1module precision_mod
2 use iso_fortran_env, only: int32, int64, real32, real64, real128
3 implicit none
4
5 integer, parameter :: prec = real64
6end module precision_mod
Listing 2: The precision module for double precision implementation.
@@
Listing 3: Extract from the module dualzn_mod
The description of the variables, functions and operators in listing 3 is as follows.
@@ 4.2 Usage of DNAOAD
The package is available for Windows and GNU/Linux in both double and quadruple precision, and it supports the gfortran and ifx Intel Fortran compilers. It can be downloaded from [30]. After uncompressing the file, the user can navigate, for instance, to the folder:
*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ (the ‘*’ depends on where the package is uncompressed on the local machine), and run some of the examples. The source code is also provided; however, in our examples, we will assume that the precompiled libraries in the GNU_Linux folder are used.
@@ 4.2.1 Simple examples
As a simple example, the program in listing 4 computes the derivatives from the 0-th up to the 5-th order of $f=\sin(z)^{\log(z^{2})}$ at $z_{0}=1.1+2.2i$. Assuming the code is saved in the folder ‘*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’ under the file example1.f90, it can be compiled in GNU/Linux as:
ifx -I../LibDualzn64 -o example1.out example1.f90 -L../LibDualzn64 -ldualzn
and in Windows as:
ifx -o example1.exe example1.f90 -I../LibDualzn64 ../LibDualzn64/libdualzn.lib
For the gfortran compiler, use:
gfortran -I../LibDualzn64 -o example1.exe example1.f90 -L../LibDualzn64 -ldualzn
Alternatively, run the provided script create_exec.sh (on GNU/Linux) or create_exec.bat (on Windows), followed by the file name where the code is saved.
@@
Listing 4: Example of derivative calculation. Since a dualzn number is an allocatable entity, it must first be initialized.
Dual numbers can also be used to differentiate Fortran code. Consider the function $f(x)=\sin(x)\exp(-x^{2})$ and the task of calculating the derivatives of $g=f(f(\cdots f(x)\cdots))$, where $f$ is nested 1,000 times. This calculation is virtually impossible symbolically, and finite differences would be inefficient and inaccurate. Listing 5 shows a program to compute these derivatives. Appendix A., presents Python and Julia versions when the funcion is nested 5 times and the order of derivation is 10. While this example is theoretical, real-world scenarios often involve multiple function compositions, vector rotations, and similar operations. For a practical demonstration, refer to [16], where kinematic quantities for the coupler point in a spherical 4R mechanism are calculated.
@@
Listing 5: Example of differentiating a function that is not given in closed form but implemented as a computer program.
@@ 4.2.2 Calculating gradients, Jacobians and Hessians
The package also includes the diff_mod module, which provides useful functions for computing gradients, Jacobians, and Hessians. The underlying theory of using dual numbers to compute these differential operators, as well as directional derivatives in general, is presented in [17]. Below an example of use of this module. This example can be found in the folder:
*/DNAOAD/GNU_Linux/ifx_case/prec_real64/examples/’.
@@
48 complex(prec), parameter :: ii = (0,1)
@@ 83 GV = gradient(fstest,q(1:3))
Listing 6: Example of using the diff_mod.
@@ 5 Conclusions
This study addresses the challenges of computing higher-order derivatives with dual numbers by introducing DNAOAD, a Fortran-based implementation designed for arbitrary-order automatic differentiation. Most existing dual-number implementations are limited to lower-order derivatives, with only one notable approach—Julia’s ForwardDiff—supporting arbitrary-order differentiation through nested dual numbers. However, no prior work has demonstrated a method for arbitrary-order differentiation with dual numbers that avoids reliance on nested structures, which are prone to significant memory consumption and stack overflow errors. DNAOAD bridges this gap by directly implementing dual numbers to arbitrary orders, eliminating the need for recursive or nested frameworks. This direct and innovative approach significantly reduces computational overhead, enabling the efficient and reliable computation of derivatives at very high orders, and providing a scalable and practical solution to the limitations of existing methods.
Through numerical experiments, DNAOAD demonstrated its capability to compute derivatives at very high orders with remarkable accuracy, though extremely high orders may require enhanced numerical precision, such as quadruple precision. The implementation also provides a considerable number of useful functions for various applications, including the calculation of gradients, Jacobians, and Hessians, as well as higher-order derivatives of complex functions. These capabilities enhance DNAOAD’s applicability to a broad range of fields, such as optimization, physics, and engineering, where precise differentiation is essential.
In comparison with existing tools, such as PyTorch and ForwardDiff—which excel in low-order differentiation but face scalability issues due to memory overheads—DNAOAD provides a scalable and efficient alternative. Its robustness was demonstrated in experiments involving deeply nested functions, highlighting its potential for tackling complex computational tasks. This implementation represents a significant advancement in practical automatic differentiation by combining theoretical rigor with computational efficiency. Future work may focus on extending DNAOAD to other programming environments and further optimizing its performance for large-scale applications.
Data availability statement
The Fortran source code, which contains the data supporting the findings of this study, is openly available on GitHub at https://github.com/fpenunuri/DNAOAD.
Appendix Appendix A. Computing higher order derivatives with PyTorch and ForwardDiff
Although the computation of higher-order derivatives can lead to significant memory consumption, the utility of the PyTorch and ForwardDiff libraries remains undeniable. Below, we present the example of Listing 5 that demonstrates a relatively small number of function compositions in both Python and Julia, respectively.
@@
Listing 7: Example of computing higher-order derivatives with PyTorch.
@@
Listing 8: Example of computing higher-order derivatives with ForwardDiff.

### v2

@@ 3.3 Dualization of elementary functions
5:  $res\leftarrow res+\binom{k}{i}A\%f(i)B\%f(k-i)$
@@
7:return resres
Algorithm 4 Product of two dual numbers.
1:$A,B$: dualzn
2:ABAB: dualzn
3:Allocate $res\%f(0:\texttt{order})$
@@
5:  $res\%f(k)\leftarrow\texttt{timesdzn}(A,B,k)$
@@ 4 The DNAOAD package
Building on the theory introduced in the previous sections, we present DNAOAD, a Fortran implementation of dual numbers of arbitrary order. In addition to enabling the computation of derivatives of arbitrary order, DNAOAD allows users to formulate numerical algorithms directly in the algebra of dual numbers, so that derivatives are propagated automatically through standard arithmetic operations and elementary functions. This section describes the structure and distribution of the package before presenting representative applications.
@@ 4.1 Elements of the package
DNAOAD is distributed in two equivalent forms. The first is a standalone source distribution intended for traditional workflows, such as building with make or integrating the source files directly into an existing Fortran code base [20]. The second is an FPM (Fortran Package Manager) distribution [18], which automatically handles compilation and linking and therefore provides a more convenient and reproducible workflow. While both distributions offer identical functionality, the FPM-based interface is recommended for new projects and for rapid integration into larger Fortran applications.
The core of the DNAOAD package is provided by the module dualzn_mod, which defines the dualzn derived type and implements the overloaded operators and intrinsic-like functions required for seamless manipulation of dual numbers. For illustration, the first 28 lines of dualzn_mod are shown in Listing 2.
@@
Listing 2: Extract from the module dualzn_mod
The description of the variables, functions and operators in listing 2 is as follows.
@@ 4.2 Usage of DNAOAD
The DNAOAD package is available for both Windows and GNU/Linux platforms in double and quadruple precision, and it supports the gfortran and Intel ifx Fortran compilers. The source code and precompiled libraries can be obtained from the project repository [20]. The distribution includes a set of example programs that illustrate the basic usage of the library.
For the standalone distribution, users may compile and run the example codes using either the provided precompiled libraries or by building the sources directly. In the examples presented here, we assume that the precompiled libraries supplied with the package are used. Detailed build instructions and platform-specific scripts are included in the distribution to simplify compilation on both GNU/Linux and Windows systems.
@@ 4.2.1 Simple examples
As a simple example, Listing 3 shows a program that computes the derivatives from zeroth to fifth order of the function $f(z)=\sin(z)^{\log(z^{2})}$ evaluated at $z_{0}=1.1+2.2\,i$. When using the standalone distribution, the program can be compiled by linking against the DNAOAD library with either the ifx or gfortran compilers, following the instructions provided with the package. For convenience, platform-specific scripts are also included to automate the compilation process.
When using the FPM (Fortran Package Manager) distribution, compilation and execution are fully managed by FPM. In this case, the example can be built and executed with a single command, for instance, fpm run ex1, which significantly simplifies the workflow.
@@
Listing 3: Example of derivative calculation. Since a dualzn number is an allocatable entity, it must first be initialized.
Dual numbers can also be used to differentiate Fortran code. Consider the function $f(x)=\sin(x)\exp(-x^{2})$ and the task of calculating the derivatives of $g=f(f(\cdots f(x)\cdots))$, where $f$ is nested 1,000 times. This calculation is virtually impossible symbolically, and finite differences would be inefficient and inaccurate. Listing 4 shows a program to compute these derivatives. Appendix A presents Python and Julia versions when the funcion is nested 5 times and the order of derivation is 10. While this example is theoretical, real-world scenarios often involve multiple function compositions, vector rotations, and similar operations. For a practical demonstration, refer to [22], where kinematic quantities for the coupler point in a spherical 4R mechanism are calculated.
@@
Listing 4: Example of differentiating a function that is not given in closed form but implemented as a computer program.
@@ 4.2.2 Calculating gradients, Jacobians and Hessians
The package also includes the diff_mod module, which provides useful functions for computing gradients, Jacobians, and Hessians. The underlying theory of using dual numbers to compute these differential operators, as well as directional derivatives in general, is presented in [21]. Below an example of use of this module.
@@
48 complex(prec), parameter :: $i$ = (0,1)
@@ 83 GV = gradient(fstest,q(1:3))
Listing 5: Example of using the diff_mod.
@@ 5 Conclusions
In this work, we introduced DNAOAD, a Fortran-based implementation of dual numbers designed to support automatic differentiation of arbitrary order. Unlike most existing dual-number approaches, which are limited to low-order derivatives or rely on recursive and nested structures, DNAOAD employs a direct, non-nested representation of dual numbers. This design avoids the severe memory growth and stack limitations commonly encountered in nested implementations and enables scalable computation of higher-order derivatives.
The numerical experiments presented in this work indicate that DNAOAD can reliably compute derivatives of substantially higher order than those typically accessible with nested dual-number implementations, without encountering stack overflows or prohibitive memory usage. While increasing the derivative order may require enhanced numerical precision—such as quadruple precision for extreme cases—the proposed approach significantly extends the range of derivative orders that can be computed in practice. In addition to higher-order derivatives, the implementation provides practical tools for computing gradients, Jacobians, Hessians, and higher-order derivatives of complex-valued functions, making DNAOAD suitable for a wide range of applications in scientific computing, physics, and engineering.
Compared with existing tools such as PyTorch and ForwardDiff, which perform well for low-order differentiation but face scalability issues due to memory overhead, DNAOAD offers a robust and efficient alternative for high-order differentiation. The results highlight the effectiveness of the proposed approach for deeply nested functions and computationally demanding problems.
Appendix A Computing higher order derivatives with PyTorch and ForwardDiff
Although the computation of higher-order derivatives can lead to significant memory consumption, the utility of the PyTorch and ForwardDiff libraries remains undeniable. Below, we present the example of Listing 4 that demonstrates a relatively small number of function compositions in both Python and Julia, respectively.
@@ Appendix A Computing higher order derivatives with PyTorch and ForwardDiff
Listing 6: Example of computing higher-order derivatives with PyTorch.
@@
Listing 7: Example of computing higher-order derivatives with ForwardDiff.